OpenStudy (pokemon23):

4(5x-2)-6x=3(4x+2) is the answer x=7

jimthompson5910 (jim_thompson5910):

yes, the answer is x=7, congrats

OpenStudy (pokemon23):

hell yeah :)

OpenStudy (mathteacher1729):

To see if the answer is 7... substitute 7 in everywhere you see an x. :) 4(5(7) -2) - 6(7) = ?? = 3(4(7)+2) If both sides are equal, then it's true. If not, then it's not true. The only way to know is to try it yourself. :)

OpenStudy (pokemon23):

i know that I check and I got a weird answer?

OpenStudy (pokemon23):

I found my mistake anyways?

OpenStudy (pokemon23):

thanks mathteacher

OpenStudy (mathteacher1729):

Break it down. Let's do 4(5(7) -2) - 6(7) first. Step 1 What did you get for 5*7 - 2 ? Step 2 What did you get when you multiplied that by 4 ? Step 3 What did you get when you subtracted 6*7 from that?

OpenStudy (pokemon23):

4(35-8)

jimthompson5910 (jim_thompson5910):

should be 4(35-2)

OpenStudy (pokemon23):

4 times 2?

jimthompson5910 (jim_thompson5910):

then you distribute to get 140-8

OpenStudy (pokemon23):

140-8-42=12=2

OpenStudy (pokemon23):

132-42=14

OpenStudy (pokemon23):

90=14

OpenStudy (pokemon23):

divide 14 to 90

jimthompson5910 (jim_thompson5910):

correct check method 4(5x-2)-6x=3(4x+2) 4(5(7)-2)-6(7)=3(4(7)+2) 4(35-2)-42 = 3(28+2) 4(33)-42 = 3(30) 132 - 42 = 90 90 = 90 ... so it works